Given isosceles triangle ABC where AB = BC = 10 and the altitude from B has length 4. Find the length of the base.

Thanks so much, I really don't understand altitude problems.

Answer by Nate(3500)

The altitude makes a right triangle:
a^2 + b^2 = c^2
a^2 + 4^2 = 10^2
a^2 + 16 = 100
a^2 = 84
a = 2*sqrt(21)
Base is twice "a":
AC = 4*sqrt(21)